# P[$\arraybackslashp#1$

### v1

@@ Abstract
Large Language Models (LLMs) have rapidly advanced and are widely adopted across diverse fields. Due to the substantial computational cost and data requirements of training from scratch, many developers choose to fine-tune or modify existing open-source models. While most adhere to open-source licenses, some falsely claim original training despite clear derivation from public models. This raises pressing concerns about intellectual property protection and highlights the need for reliable methods to verify model provenance. In this paper, we propose GhostSpec, a lightweight yet effective method for verifying LLM lineage without access to training data or modification of model behavior. Our approach constructs compact and robust fingerprints by applying singular value decomposition (SVD) to invariant products of internal attention weight matrices, effectively capturing the structural identity of a model. Unlike watermarking or output-based methods, GhostSpec is fully data-free, non-invasive, and computationally efficient. It demonstrates strong robustness to sequential fine-tuning, pruning, block expansion, and even adversarial transformations. Extensive experiments show that GhostSpec can reliably trace the lineage of transformed models with minimal overhead. By offering a practical solution for model verification and reuse tracking, our method contributes to the protection of intellectual property and fosters a transparent, trustworthy ecosystem for large-scale language models.
@@ 1 Introduction
Large Language Models (LLMs) have rapidly advanced and become foundational infrastructure for a wide range of applications across research and industry (achiam2023gpt; yang2025qwen3; dubey2024llama). However, the high computational and data costs of training LLMs from scratch (workshop2022bloom) have led many developers to build their own models by modifying existing open-source LLMs through techniques such as instruction fine-tuning, continued pre-training, model merging, and compression (yang2024model; zhu2024survey; tang2025spindlekv). While most developers comply with open-source licenses, there have been instances of falsely claiming to have trained models “from scratch” when they are in fact repackaged or fine-tuned versions of public models (e.g., Llama3-V and MiniCPM-Llama3-V 2.5) (yao2024minicpm). This trend raises serious concerns regarding model plagiarism and intellectual property violations, highlighting the urgent need for reliable tools to verify model lineage.
@@
In this work, we propose GhostSpec, a simple, data-free, and robust white-box method for verifying the lineage of LLMs. Our key insight is that the spectral structure of weight matrices encodes intrinsic information about a model’s origin, remaining stable under various modifications. Specifically, we apply singular value decomposition (SVD) to internal matrix products within the attention mechanism, including the query-key and value-output weight products. From the resulting spectra, we select the most significant singular values based on effective rank to compute similarity. To handle architectural variations in depth, such as those resulting from layer pruning or expansion, we develop the Penalty-based Optimal Spectral Alignment (POSA) algorithm that finds the best layer-wise correspondence between models. This yields a quantitative similarity score robust to differences in depth and architectural variations. Unlike black-box or representation-based white-box methods, GhostSpec is data-independent, requires no model modification, and has minimal computational cost.
@@
  • We propose GhostSpec, a lightweight white-box method for verifying LLM lineage directly from model weights. It requires no training data or architectural changes, and remains robust to fine-tuning and pruning—offering a practical solution for provenance verification and IP protection in open-source LLMs.
  • We introduce spectral fingerprints based on invariant matrix products within the attention mechanism, which are intrinsically robust to functionality-preserving scaling and permutation transformations. Additionally, we develop the Penalty-based Optimal Spectral Alignment (POSA) algorithm to enable robust comparison across models with varying depths and architectures.
  • We conduct extensive experiments demonstrating that GhostSpec effectively distinguishes derivative models from independently trained ones, even under challenging modification scenarios, validating its utility and reliability.
@@ 2.2 White-Box Identification
This line of work focuses on static, data-free analysis of model weights. Prior methods include visualizing invariant structural features (zeng2024huref) or analyzing layer-wise statistics (yoon2025intrinsic). Our proposed GhostSpec method falls into this category, capturing deeper structural information by leveraging the full singular value spectrum of invariant matrix products. The Penalty-based Optimal Spectral Alignment(POSA) algorithm is further introduced to robustly trace model ancestry under various transformations.
@@ 3 Preliminaries
staats2024small employ Random Matrix Theory (RMT) to analyze the singular value spectrum of weight matrices in pre-trained language models. Their analysis reveals significant deviations from the Marchenko–Pastur distribution expected from randomly initialized weight matrices, especially in the region of large singular values. These outliers correspond to dominant directions in the weight space and are strongly associated with the model’s learned representations. Removing them leads to a substantial increase in perplexity, indicating that they form a stable backbone of the model’s identity.
Fine-tuning, in contrast to training from scratch, refines rather than rebuilds the model’s internal structure. staats2024small demonstrate that fine-tuning primarily affects directions associated with small singular values. A key asymmetry is observed: pruning small singular values after fine-tuning leads to significantly greater performance degradation than pruning them before fine-tuning. This indicates that fine-tuning updates are concentrated in low-magnitude spectral components.
Together, these findings suggest spectral stability: large singular values, encoding foundational pre-trained knowledge, remain stable throughout adaptation and anchor the model’s global behavior. In contrast, fine-tuning introduces targeted, low-rank modifications reflected in small singular values. Based on this observation, we hypothesize that the large singular value spectrum of attention matrices serves as a robust fingerprint of a model, largely invariant to fine-tuning.
@@ 3.1 Experimental Design
Our objective is to quantify the spectral distance between the base model and each model in these groups, expecting minimal distance for fine-tuned variants and a significantly larger distance for unrelated models.
@@ 3.2 Quantifying Spectral Similarity
$\begin{split}d_{p}^{(i)}(A,B)=\frac{1}{r_{p}^{(i)}}\left\lVert\right.&\text{norm}(\text{trunc}(\mathbf{s}_{p,A}^{(i)}))\\ &\left.-\text{norm}(\text{trunc}(\mathbf{s}_{p,B}^{(i)}))\right\rVert_{2}^{2},\end{split}$ (1)
where $\text{trunc}(\cdot)$ denotes truncation to rank $r_{p}^{(i)}$ and $\text{norm}(\cdot)$ denotes min-max normalization.
@@ 3.3 Empirical Validation
As shown in Figure 2, fine-tuned models exhibit a low average spectral MSE from their base models, indicating near-identical structural characteristics. In contrast, unrelated models yield significantly higher distances, highlighting their structural dissimilarity.
These results validate that the singular value spectrum is a stable and intrinsic property of a model, largely preserved through fine-tuning. This motivates the design of a robust fingerprinting framework capable of verifying model lineage under various transformations.
@@ 4 Methodology
In practice, model weights may undergo transformation attacks such as permutation or scaling, which preserve the model’s functionality while significantly altering the raw weight distribution (zhang2024reef). Direct comparison of the singular values of individual attention matrices $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ is therefore highly vulnerable to such obfuscation techniques. To address this, we propose a robust spectral fingerprinting methodology built on two core principles: (1) an invariant fingerprint derived from composite matrix products, and (2) two complementary similarity metrics, Ghost-mse and Ghost-corr, designed to jointly capture fine-grained variations and macroscopic structural properties.
@@ 4.1 The GhostSpec Fingerprint
We define two invariant matrices per layer whose singular value spectra are resilient to adversarial attacks (see Appendix A for proof):
@@ 4.2 Similarity Metrics
Given the GhostSpec fingerprints of two models, $A$ and $B$, we introduce two complementary metrics to quantify their similarity.
@@
This metric performs a detailed, layer-by-layer comparison of the singular value vectors to populate an aggregate distance matrix $D_{\text{avg}}\in\mathbb{R}^{N\times M}$. Each entry $(D_{\text{avg}})_{ij}$ represents the average spectral distance between layer $i$ of model A and layer $j$ of model B, computed as the mean of the Mean Squared Error (MSE) over the invariant components:
@@
To handle models with different depths ($N\neq M$), we apply our Penalty-based Optimal Spectral Alignment (POSA) algorithm 1 to find the minimum-cost alignment path through $D_{\text{avg}}$. The raw similarity score, $d_{\text{path}}$, is the average MSE along this optimal path. Finally, we convert this distance into a normalized similarity score using an inverted Sigmoid transformation:
@@
This metric captures the macroscopic relationship between models by comparing the overall trend of their spectral properties. For each layer $i$ and component $p$, we compute a single scalar value, $\mu_{p,M}^{(i)}$, which is the mean of the top-$K$ normalized singular values.
This process yields two trend sequences for each model: $\boldsymbol{\mu}_{qk,M}$ and $\boldsymbol{\mu}_{vo,M}$. After aligning the sequences to handle differing depths, we concatenate them and compute the final similarity score using the distance correlation coefficient:
$\begin{split}\mathrm{Sim}_{\text{Corr}}(A,B)=\mathrm{dCor}\big(&[\boldsymbol{\mu}_{qk,A};\boldsymbol{\mu}_{vo,A}],\\ &[\boldsymbol{\mu}_{qk,B};\boldsymbol{\mu}_{vo,B}]\big).\end{split}$ (7)
This score measures the correlation of the models’ high-level spectral evolution, providing a computationally efficient indicator of shared lineage.
@@ 5.1 Experimental Setup
$\tau^{*}=\operatorname*{arg,max}_{\tau\in\mathcal{S}}\text{F1}(\hat{y}_{i}(\tau),y_{i}).$ (8)
@@ 5.2 Main Results
The overall classification performance of each method, measured by its maximum F1-score on our dataset, is summarized in Figure 4. Our two proposed variants, GhostSpec-corr (F1 = 0.9730) and GhostSpec-mse (F1 = 0.9867), achieve a clear lead, outperforming all data-aware and data-free baselines. Table  1 presents a selection of the detailed similarity scores, color-coded using these optimal thresholds to visualize the classification results. The following sections detail the performance across specific modification scenarios.
Robustness to Fine-Tuning.
Fine-tuning is a common source of model derivation. We evaluate similarity between Llama-2-7B and several of its fine-tuned variants: Llama-2-7B-Chat-HF, Vicuna-7B-v1.5, Xwin-LM-7B-V0.2, and Llemma-7B, etc. Despite differences in training data and objectives, GhostSpec consistently produces high similarity scores.
Robustness to Model Pruning.
We assess GhostSpec’s robustness under model pruning. For structured pruning, we consider methods that remove layers based on importance, such as ShortGPT (men2024shortgpt) and Sheared Llama (xia2023sheared). For unstructured pruning, we evaluate representative importance-based methods that prune weights according to their significance, including Wanda (sun2023simple) and SparseGPT (frantar2023sparsegpt). Prior work (ma2025model) identifies critical sparsity thresholds near 35% (structured) and 60% (unstructured), beyond which model behavior and redundancy shift significantly. We thus evaluate sparsity levels at 30%, 50%, and 70%, finding that GhostSpec consistently and accurately recovers the source model across all settings.
Robustness to Model Merging.
Model merging combines weights from multiple source models to create a new model. We evaluate GhostSpec on various merging approaches, including direct weight averaging, methods that reduce parameter interference through pruning and sign conflict resolution (yadav2023ties), as well as techniques that merge models by fusing weights or aligning output distributions (goddard-etal-2024-arcees; wan2024knowledge). Across all these methods, GhostSpec consistently detects strong similarity between the merged models and their original parents.
Robustness to Model Upcycling.
Model upcycling expands model capacity by introducing new components, such as additional layers or sparsely activated expert modules, including techniques that convert dense LLMs into sparse MoEs through Parameter-Efficient Sparsity Crafting (PESC) (wu2024parameter). Despite architectural modifications, GhostSpec consistently yields high similarity scores with the original base models.
Robustness to Permutation and Scaling Transformations.
To evaluate invariance under adversarial weight manipulations, we apply random permutations to the hidden dimensions of MLP weight matrices and introduce random scaling factors to the attention layers. Despite these aggressive transformations, GhostSpec consistently recognizes the original and perturbed models as similar, validating its theoretical invariance guarantees.
Robustness to Mistral-based Derivatives.
To further validate the architectural generalizability of GhostSpec, we extend our evaluation to the Mistral model family beyond the Llama series. This evaluation covers a diverse set of Mistral-based derivatives, including instruction-tuned models, DPO-optimized variants (nous2024hermes), compressed versions using pruning, or sparsification, as well as models with structural changes such as merging, interpolation, and MoE-style expansion. Despite the diversity of these modifications, GhostSpec consistently maintains accurate and reliable lineage detection across all cases, demonstrating its strong robustness and cross-architecture transferability.
Discrimination Against Unrelated Models.
To evaluate the specificity of GhostSpec and minimize false positives, we tested its performance on architecturally unrelated models. We computed similarity scores between our base models (Llama-2-7b and Mistral-7B) and a diverse set of independently trained LLMs, including pythia-6.9b, opt-6.7b, among others(MosaicML2023Introducing; guo2024deepseek; zhang2022opt). GhostSpec consistently assigned near-zero similarity scores, reliably distinguishing true derivatives from unrelated models and demonstrating strong discriminative capability.
Figure 5: Pairwise structural similarity matrix of prominent open-source models computed using GhostSpec-mse. Higher scores indicate greater similarity.

### v2

@@ Abstract
Large Language Models (LLMs) are widely adopted, but their high training cost leads many developers to fine-tune existing open-source models. While most adhere to open-source licenses, some falsely claim original training despite clear derivation from public models, raising pressing concerns about intellectual property protection and the need to verify model provenance. In this paper, we propose GhostSpec, a lightweight yet effective method for verifying LLM lineage without access to training data or modification of model behavior. Our approach constructs compact and robust fingerprints by applying singular value decomposition (SVD) to invariant products of internal attention weight matrices. Unlike watermarking or output-based methods, GhostSpec is fully data-free, non-invasive, and computationally efficient. Extensive experiments show it is robust to fine-tuning, pruning, expansion, and adversarial transformations, reliably tracing lineage with minimal overhead. By offering a practical solution for model verification, our method contributes to intellectual property protection and fosters a transparent, trustworthy LLM ecosystem. Our code is available at https://github.com/DX0369/GhostSpec.
Extended versionhttp://arxiv.org/abs/2511.06390
@@ 1 Introduction
LLMs have quickly become essential for various applications in research and industry (achiam2023gpt; yang2025qwen3; wang2025parameters; zhang2025segment; poon2025online). Due to the high cost of training LLMs from scratch (workshop2022bloom), many developers modify open-source LLMs via fine-tuning, continued pre-training, merging, and compression (yang2024model; zhu2024survey; tang2025covipal; yang2025xquant; li2023enhancing; hu2025songsong). While most developers comply with open-source licenses, there have been instances of falsely claiming to have trained models “from scratch” when they are in fact repackaged or fine-tuned versions of public models (e.g., Llama3-V and MiniCPM-Llama3-V 2.5) (yao2024minicpm). It is crucial to distinguish such intellectual property violations, which often break attribution requirements, from legitimate, licensed fine-tuning. This raises concerns about plagiarism and intellectual property violations, emphasizing the need for tools to verify model lineage.
@@
We propose GhostSpec, a simple, data-free, and robust white-box method for verifying LLM lineage. GhostSpec is, by design, a white-box method targeting the open-weight model ecosystem where weights are accessible. Our key insight is that the spectral structure of weight matrices encodes intrinsic information about a model’s origin, remaining stable under various modifications. Specifically, we apply singular value decomposition (SVD) to internal matrix products within the attention mechanism, including the query-key and value-output weight products. From the resulting spectra, we select the most significant singular values based on effective rank to compute similarity. To handle architectural variations in depth, such as those resulting from layer pruning or expansion, we develop the Penalty-based Optimal Spectral Alignment (POSA) algorithm that finds the best layer-wise correspondence between models. This yields a quantitative similarity score robust to differences in depth and architectural variations. Unlike black-box or representation-based white-box methods, GhostSpec is data-independent, requires no model modification, and has minimal computational cost.
@@
  • We propose GhostSpec, a lightweight white-box method for verifying LLM lineage from model weights, requiring no training data or architectural changes, offering a practical solution for provenance verification and IP protection in open-source LLMs.
  • We introduce spectral fingerprints based on invariant matrix products within the attention mechanism, robust to scaling and permutation transformations, and develop the POSA algorithm for comparing models with varying depths and architectures.
  • We demonstrate through experiments that GhostSpec reliably distinguishes derivative models from independently trained ones, even under challenging modifications.
@@ 2.2 White-Box Identification
This line of work focuses on static, data-free analysis of model weights. Prior methods include visualizing invariant structural features (zeng2024huref) or analyzing layer-wise statistics (yoon2025intrinsic). Our proposed GhostSpec method falls into this category, capturing deeper structural information by leveraging the full singular value spectrum of invariant matrix products. The POSA algorithm is further introduced to robustly trace model ancestry under various transformations.
@@ 3 Preliminaries
staats2024small use Random Matrix Theory (RMT) to analyze the singular value spectrum of pre-trained language models. Their analysis reveals deviations from the Marchenko–Pastur distribution, especially in large singular values. These outliers correspond to dominant directions in the weight space and are strongly associated with the model’s learned representations. Removing them significantly increases perplexity, showing they form a stable backbone of the model’s identity.
Fine-tuning, in contrast to training from scratch, refines rather than rebuilds the model’s internal structure. staats2024small demonstrates that fine-tuning primarily affects directions associated with small singular values. A key asymmetry is observed: pruning small singular values after fine-tuning leads to significantly greater performance degradation than pruning them before fine-tuning. This indicates that fine-tuning updates are concentrated in low-magnitude spectral components.
Together, these findings suggest spectral stability: large singular values, encoding foundational pre-trained knowledge, remain stable throughout adaptation and anchor the model’s global behavior. In contrast, fine-tuning introduces targeted, low-rank modifications reflected in small singular values. We hypothesize that the large singular value spectrum of attention matrices serves as a robust, fine-tuning-invariant fingerprint.
@@ 3.1 Experimental Design
Our objective is to quantify the spectral distance between the base model and each model in these groups.
@@ 3.2 Quantifying Spectral Similarity
$\begin{split}d_{p}^{(i)}(A,B)=\frac{1}{r_{p}^{(i)}}\left\lVert\right.&\text{norm}(\text{trunc}(\mathbf{s}_{p,A}^{(i)}))\\ &\left.-\text{norm}(\text{trunc}(\mathbf{s}_{p,B}^{(i)}))\right\rVert_{2}^{2}.\end{split}$ (1)
@@ 3.3 Empirical Validation
As shown in Figure 2, fine-tuned models have low spectral MSE with their base models, indicating similar structures. In contrast, unrelated models show much higher distances, highlighting structural dissimilarity.
These results confirm that the singular value spectrum is a stable, intrinsic property, mostly preserved through fine-tuning, motivating the design of a robust fingerprinting framework for model lineage verification under transformations.
@@ 4 Methodology
In practice, model weights may undergo transformation attacks such as permutation or scaling, which preserve the model’s functionality while significantly altering the raw weight distribution (zhang2024reef). Directly comparing the singular values of individual attention matrices $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$ is therefore highly vulnerable to such obfuscation techniques. To address this, we propose constructing an invariant fingerprint derived from composite matrix products, which are resistant to these transformation attacks. Additionally, we introduce two complementary similarity metrics, GhostSpec-mse and GhostSpec-corr, designed to capture both fine-grained variations and macroscopic structural properties. Together, these components form a robust spectral fingerprinting methodology, as shown in Figure 1.
@@ 4.1 The GhostSpec Fingerprint
We define two invariant matrices per layer whose singular value spectra are resilient to functionality-preserving transformations like permutation or scaling:
@@
These transformations alter the weight distribution without changing the model’s output by modifying the weights in pairs or scaling them uniformly. Since the relative relationships between the weights are preserved, the model’s core functionality and input-output behavior remain unchanged.
@@ 4.2 Similarity Metrics
Given the GhostSpec fingerprints of two models, $A$ (with $N$ layers) and $B$ (with $M$ layers), we introduce two complementary metrics: GhostSpec-mse, which provides a fine-grained, layer-by-layer comparison to measure structural correspondence, and GhostSpec-corr, a lightweight metric that captures the overall trend of spectral properties across layers. Used together, these metrics offer a comprehensive and reliable assessment of model lineage.
@@
This metric performs a detailed, layer-by-layer comparison of the singular value vectors to populate an aggregate distance matrix $D_{\text{avg}}\in\mathbb{R}^{N\times M}$. Each entry $(D_{\text{avg}})_{ij}$ represents the average spectral distance between layer $i$ of model A and layer $j$ of model B, computed as the mean of MSE over the invariant components:
@@
To handle models with different depths ($N\neq M$), we apply our POSA algorithm, as shown in Algorithm 1, to find the minimum-cost alignment path through $D_{\text{avg}}$. The raw similarity score, $d_{\text{path}}$, is the average MSE along this optimal path. Finally, we convert this distance into a normalized similarity score using an inverted Sigmoid transformation:
@@
As illustrated in Figure 3, models with shared lineage exhibit highly correlated spectral trends, while unrelated models show divergent patterns. Based on this observation, we propose the GhostSpec-corr metric, which quantifies model similarity by capturing the overall trend of spectral properties.
First, we generate the trend sequences. For each layer $i$ and component $p$, we compute a scalar value, $\mu_{p,M}^{(i)}$, defined as the mean of the top-$K$ normalized singular values. Here, $K$ is dynamically determined based on the effective rank of each singular value spectrum, consistent with the truncation method used in GhostSpec-mse. This process produces two trend sequences for each model: $\boldsymbol{\mu}_{qk,M}$ and $\boldsymbol{\mu}_{vo,M}$.
Second, we align the sequences. Since these sequences may differ in length, direct comparison is not feasible. To address this, we apply a dynamic sequence alignment algorithm (similar to the POSA algorithm introduced previously) to match the sequence lengths. This step generates new sequences of equal length ($\boldsymbol{\mu}^{\prime}_{qk,A}$, $\boldsymbol{\mu}^{\prime}_{qk,B}$, etc.).
Finally, we compute the similarity. We concatenate the aligned sequences for each model and calculate the final similarity score using the distance correlation coefficient:
$\begin{split}\mathrm{Sim}_{\text{Corr}}(A,B)=\mathrm{dCor}\big(&[\boldsymbol{\mu}^{\prime}_{qk,A};\boldsymbol{\mu}^{\prime}_{vo,A}],\\ &[\boldsymbol{\mu}^{\prime}_{qk,B};\boldsymbol{\mu}^{\prime}_{vo,B}]\big).\end{split}$ (7)
This score quantifies the correlation of the models’ high-level spectral evolution, providing a computationally efficient indicator of shared lineage.
@@ 5.1 Experimental Setup
  • Fine-Tuning. Fine-tuning is a common source of model derivation. We evaluate similarity between Llama-2-7B and several of its fine-tuned variants: Llama-2-7B-Chat-HF, Vicuna-7B-v1.5 and Llemma-7B, etc. Mistral-7B and several of its fine-tuned variants, including instruction-tuned models, DPO-optimized variants (nous2024hermes).
  • Model Pruning. For structured pruning, we consider methods that remove layers based on importance, such as ShortGPT (men2024shortgpt) and Sheared Llama (xia2023sheared). For unstructured pruning, we evaluate representative importance-based methods that prune weights according to their significance, including Wanda (sun2023simple) and SparseGPT (frantar2023sparsegpt). ma2025model identifies critical sparsity thresholds near 35% (structured) and 60% (unstructured), beyond which model performance rapidly deteriorates. We thus evaluate sparsity levels at 30%, 50%, and 70%.
  • Model Merging. Model merging combines weights from multiple source models to create a new model. We evaluate GhostSpec on various merging approaches, including direct weight averaging, methods that reduce parameter interference through pruning and sign conflict resolution (yadav2023ties), as well as techniques that merge models by fusing weights or aligning output distributions (goddard-etal-2024-arcees; wan2024knowledge).
  • Model Upcycling. Model upcycling expands model capacity by introducing new components, such as additional layers or sparsely activated expert modules, including techniques that convert dense LLMs into sparse MoEs through Parameter-Efficient Sparsity Crafting (PESC) (wu2024parameter).
  • Permutation and Scaling Transformations. To evaluate invariance under adversarial weight manipulations, we apply random permutations to the hidden dimensions of MLP weight matrices. For the attention layers, we apply functionality-preserving similarity transforms, which combine both random rotation and scaling, to the attention projection matrices.
  • Unrelated Models. We computed similarity scores between our base models (Llama-2-7b and Mistral-7B) and a diverse set of independently trained LLMs, including pythia-6.9b, opt-6.7b, among others(MosaicML2023Introducing; guo2024deepseek; zhang2022opt).
@@
$\tau^{*}=\operatorname*{argmax}_{\tau\in\mathcal{S}}\text{F1}(\hat{y}_{i}(\tau),y_{i}).$ (8)
where:
  • $y_{i}$ is the true label, which is 1 for related pairs (based on ground truth) and 0 for unrelated pairs.
  • $\hat{y}_{i}(\tau)$ is the predicted label, which is 1 if the similarity score exceeds the threshold $\tau$, and 0 otherwise.
@@ 5.2 Main Results
The overall classification performance of each method, measured by its maximum F1-score on our dataset, is summarized in Figure 4. Our two proposed variants, GhostSpec-corr (F1 = 0.9730) and GhostSpec-mse (F1 = 0.9867), achieve a clear lead, outperforming all data-aware and data-free baselines. Table 1 presents a selection of illustrative examples from our 63-pair dataset, showing their detailed similarity scores. These scores are color-coded using the method-specific optimal thresholds, providing an intuitive visualization of the final classification results
GhostSpec demonstrated robust performance across various model transformations. It consistently produces high similarity scores for fine-tuned variants, regardless of training data or objectives. This robustness extends to structural modifications, including aggressive pruning (up to 70% sparsity), where it accurately recovers source models. It also detects strong similarity in merged and upcycled models, and remains resilient to adversarial transformations like permutation and scaling. Overall, GhostSpec reliably distinguishes true derivatives from unrelated models, showcasing strong discriminative power.
Figure 5: Pairwise structural similarity matrix of prominent open-source models computed using GhostSpec-mse. The heatmap visualizes the genealogical relationships between prominent open-source models. Higher scores indicate greater similarity.